OpenStudy (anonymous):

What is the exact value of the expression the square root 72. − the square root of 8. + the square root of 128.? Simplify if possible. 8the square root of 2. 12the square root of 2. 8the square root of 3. 12the square root of 3.

OpenStudy (anonymous):

\[\sqrt{72}-\sqrt{8}+\sqrt{128}\]?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

you are going to get a bunch of terms with \(\sqrt{2}\) in them, so ignore the last two choices

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

one by one \[\sqrt{72}=\sqrt{36\times 2}=\sqrt{36}\sqrt{2}=6\sqrt{2}\]

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

clear so b ?

OpenStudy (anonymous):

idk we are not nearly done that was just one term !

OpenStudy (anonymous):

what about \(\sqrt{8}\) ?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

it is clear that \(8=4\times 2\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which gives you \(\sqrt{8}=\sqrt{4}\sqrt{2}=2\sqrt{2}\)

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

how about \(128\) ?

OpenStudy (anonymous):

i have no idea if it is A or B, i know it is not C or D because they both have \(\sqrt3\) in them

OpenStudy (anonymous):

we still have to figure out how to deal with \(\sqrt{128}\)

OpenStudy (anonymous):

what is half of 128 ?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

ok good what is the square root of 64 ?

OpenStudy (anonymous):

8

OpenStudy (primeralph):